OpenStudy (anonymous):

Find the vertex of the parabola by completing the square. x^2 - 6x + 8 = y A. (–3,– 6) B. (3,–1) C. (–3,1) D. (–1, 3) A, B, C, or D? Please explain :)

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{2a}\) which in your case is \(-\frac{-6}{2\times 1}=3\) second coordinate is what you get when your replace \(x\) by \(3\)

OpenStudy (anonymous):

Thanks for the explanation you got any idea on the answer? What's your point of view on that?

OpenStudy (raden):

3^2-6(3)+8 = ...